# Meta-Reasoner: Dynamic Guidance for Optimized Inference-time Reasoning in Large Language Models

### v3

@@
Abstract
Large Language Models (LLMs) increasingly rely on prolonged reasoning chains to solve complex tasks. However, this trial-and-error approach often leads to high computational overhead and error propagation, where early mistakes can derail subsequent steps. To address these issues, we introduce Meta-Reasoner, a framework that dynamically optimizes inference-time reasoning by enabling LLMs to “think about how to think.” Drawing inspiration from human meta-cognition and dual-process theory, Meta-Reasoner operates as a strategic advisor, decoupling high-level guidance from step-by-step generation. It employs contextual multi-armed bandits to iteratively evaluate reasoning progress and select optimal strategies (e.g., backtrack, clarify ambiguity, restart from scratch, or propose alternative approaches), and reallocates computational resources toward the most promising paths. Our evaluations on mathematical reasoning and puzzles highlight the potential of dynamic reasoning chains to overcome inherent challenges in the LLM reasoning process and also show promise in broader applications, offering a scalable and adaptable solution for reasoning-intensive tasks.
Meta-Reasoner: Dynamic Guidance for Optimized Inference-time
Reasoning in Large Language Models
Yuan Sui1, Yufei He1, Tri Cao1, Simeng Han2, Yulin Chen1, Bryan Hooi1 1 National University of Singapore 2 Yale University yuan.sui@u.nus.edu, bhooi@comp.nus.edu.sg
1 Introduction
@@
Recent advances in o1-like and r1-like [^1] O1 (from OpenAI) and R1 (from Deepseek) are two advanced LLM models designed for reasoning. reasoning have enabled large language models (LLMs) to achieve remarkable performance on complex tasks such as mathematics Patel et al. (2024); Lightman et al. (2023), science Rein et al. (2023), and logical puzzles Lei et al. (2024); Yao et al. (2023). By simulating multi-step, human-like deliberation Yao et al. (2024), these methods allow LLMs to decompose problems into smaller subproblems, test hypotheses, reflect on intermediate results, and iteratively refine their solutions. This extended reasoning process enables systematic exploration of ideas, verification of partial conclusions, and progressive improvement before producing a final answer. Such capabilities are particularly valuable in domains demanding rigorous logical reasoning Chenghao Yang (2024).
Despite these advances, o1/r1-like reasoning remains fundamentally challenged by its trial-and-error nature: models generate numerous candidate reasoning paths, discard flawed ones, and gradually converge on solutions. While this flexibility facilitates exploration of diverse strategies, it often incurs substantial computational overhead Snell et al. (2024); Manvi et al. (2024) and is vulnerable to error propagation, where early mistakes accumulate and compromise subsequent steps Lei et al. (2024); Yao et al. (2023); Gandhi et al. (2024). Some iterative methods incorporate partial revision or backtracking Gandhi et al. (2024); Li et al. (2025a), but these approaches tend to be ad-hoc and limited to correcting errors within a narrow reasoning window. Crucially, they lack a systematic mechanism to assess whether an entire reasoning trajectory remains promising or should be abandoned.
As a result, LLMs risk becoming “stuck” on unproductive reasoning paths, wasting valuable computational resources without recognizing when a strategic pivot is necessary. A critical challenge, therefore, is to enable LLMs to manage their reasoning budget more effectively—prioritizing promising directions while adapting or discarding ineffective strategies during inference time.
To address this challenge, we propose Meta-Reasoner, a specialized meta-reasoning module that operates alongside the LLM to enhance its reasoning capabilities. Acting as a high-level advisor, the meta-reasoner dynamically evaluates the reasoning process and provides strategic guidance or redirection when progress stalls. Unlike the LLM, which focuses on detailed stepwise generation, the meta-reasoner maintains a global perspective, assessing overall progress and strategy from a high level. Meta-Reasoner operates in iterative rounds: First, the LLM generates partial chain-of-thought reasoning chains and a concise “progress report” summarizing its current state. Meta-Reasoner then reviews this report and offers high-level feedback - such as restarting reasoning with a different approach, refining existing ideas, or focusing on specific subproblems. This setup allows Meta-Reasoner to concentrate on overall strategy rather than getting involved in the granular details of the LLM’s reasoning. Overall, Meta-Reasoner helps prevent the LLM from getting stuck or spending resources on unproductive lines of inquiry during the inference time.
Overall, the main contributions are as follows:
@@
  • We propose a meta-reasoning module that provides high-level oversight and guidance to the LLM’s reasoning, helping to prevent it from getting stuck on unproductive solution paths during inference.
@@
  • We empirically show that our Meta-Reasoner improves accuracy and efficiency on challenging math and scientific reasoning benchmarks (e.g., Game of 24, TheoremQA, and SciBench) compared to existing strong baselines. Our results highlight the framework’s potential as a new scalable solution to inference-time reasoning.
2 Related Works
@@
Complex Reasoning in LLMs
@@
The introduction of CoT reasoning has transformed how LLMs tackle complex problems by decomposing tasks into intermediate steps Lee et al. (2025). Recent models such as OpenAI’s o1 and o3, and Deepseek’s r1, have achieved state-of-the-art performance across diverse domains by leveraging CoT-like reasoning Manvi et al. (2024); Li et al. (2025a); Kudo et al. (2024); Sui et al. (2024b). However, the inherent sequential dependency in CoT limits robustness: errors in early steps can propagate and degrade overall performance Snell et al. (2024). Moreover, when confronted with complex reasoning tasks, CoT methods may become trapped in infinite reasoning loops Lee et al. (2025); Sui et al. (2024a, c). To address these challenges, we propose Meta-Reasoner, a meta-reasoner that monitors and adapts the reasoning strategy based on the progress of CoT. By integrating mechanisms such as backtracking and self-verification, Meta-Reasoner provides a holistic perspective that evaluates the overall reasoning trajectory to prevent stagnation and reduce wasted computation on unproductive lines of thought.
Backtracking and Self-Verification
@@
To mitigate the limitations of CoT-like reasoning, recent methods have explored backtracking and self-verification techniques Yao et al. (2023); Besta et al. (2023); Gandhi et al. (2024). For example, Weng et al. (2023) demonstrate that incorporating a self-verification stage—where the model re-examines its conclusions using the generated chain of thought—significantly improves performance by detecting errors early. Similarly, Ling et al. (2023) propose generating multiple candidate reasoning chains alongside a verifier mechanism that identifies and backtracks on erroneous steps. These approaches extend beyond post-hoc validation by enabling dynamic strategy adjustments during inference Lightman et al. (2023), thereby limiting error propagation in lengthy reasoning chains and mitigating infinite reasoning loops. Building on these efforts, our Meta-Reasoner framework employs instructions to (1) restart from scratch with alternative strategies, (2) backtrack to the error point, and (3) continue with targeted suggestions. Further details on this strategy are provided in §4.3.
Meta-Cognition & Dual-Process Systems
@@
From a cognitive science perspective, meta-cognition involves higher-order processes that allow individuals to monitor, evaluate, and adjust their cognitive strategies Gao et al. (2024); Yoran et al. (2024). This reflective thinking—often characterized as System 2 in dual-process theories Havrilla et al. (2024)—is vital for tasks requiring careful deliberation and error correction Didolkar et al. (2024). Drawing on these insights, our Meta-Reasoner framework can be viewed as analogous to dual-process systems: the LLM generates CoT steps akin to System 1, while the Meta-Reasoner provides high-level strategic oversight, analogous to System 2, guiding or redirecting reasoning as needed. This separation of responsibilities balances efficiency with robust problem-solving, allowing the LLM to handle routine inferences and the Meta-Reasoner to intervene for strategic adjustments.
3 Preliminary
In complex reasoning tasks, a key challenge is deciding which strategy to use from multiple valid options. This problem can be naturally framed as a contextual multi-armed bandit (MAB) problem, a well-studied framework for making decisions that balance trying new options (exploration) and using what is already known to work (exploitation).
@@ 3 Preliminary
Imagine an agent faced with several strategies (called arms), and at each step, it observes some information about the current situation (called the context). Based on this context, the agent picks one strategy to apply and then receives feedback (a reward) indicating how well that strategy performed. The goal is to choose strategies over time to maximize the total reward.
@@
$R(T)=\sum_{t=1}^{T}r(s_{t},x_{t}).$ (1)
A popular algorithm for this problem is LinUCB Li et al. (2012), which assumes the expected reward is approximately a linear function of the context. Specifically, for each arm $s$, there is an unknown parameter vector $\theta_{s}$ such that:
$\mathbb{E}[r(s,x_{t})]\approx x_{t}^{\top}\theta_{s}.$ (2)
@@
$s_{t}=\arg\max_{s\in\mathcal{S}}\left[x_{t}^{\top}\hat{\theta}_{s}+c\sqrt{x_{t}^{\top}A_{s}^{-1}x_{t}}\right],$ (3)
@@
By using context to guide its choices, LinUCB adapts its strategy selection to the current situation, which aligns with our goal in Meta-Reasoner: to dynamically choose the most effective reasoning strategy during inference time.
4 Methods
Motivated by the intuition that LLMs should concentrate their computational efforts on more promising reasoning paths during inference time, we explore two key research questions in this paper: (1) How can language models dynamically allocate resources during inference to optimize reasoning and planning?; (2) What architectural design enables an effective separation between the reasoning process within the LLM and the meta-level guidance that oversees it?
To address these questions, we propose a novel framework, Meta-Reasoner, which endows LLMs with the ability to “think about how to think”. Our framework supervises the reasoning process and dynamically guides the model to focus on more promising reasoning trajectories during inference time. Furthermore, Meta-Reasoner mitigates the limitations of conventional sequential reasoning, which may get stuck in suboptimal paths. We propose a “high-order" reasoning mechanism to balance exploration and exploitation using a Multi-Armed Bandit (MAB) algorithm.
@@ 4 Methods
Figure 1: An illustration of the Meta-Reasoner workflow. In each round, the LLM produces a new reasoning step to extend its CoT reasoning. The CoT is then summarized into a progress report, which provides context for the meta-reasoner. Then meta-reasoner uses a contextual multi-armed bandit (either using a fixed contextual bandit or dynamic contextual bandit) to choose a guidance strategy. The selected strategy then guides the next reasoning step generation, to enable strategic redirection, error correction, and resource optimization. A reward is then computed from the progress report and used to update the bandit algorithm. The process repeats until the task is complete or the maximum number of rounds is reached.
The meta-reasoning framework operates iteratively as illustrated in Figure 1. At each round $t$, the reasoning process comprises three steps: (1) CoT generation by the LLM, (2) Progress Reporting to summarize the reasoning progress so far (i.e., this is partly for efficiency, and partly to help the meta-reasoner focus on its main goal of “advising” rather than being distracted by the details in the CoT), and (3) Strategy Generation by the meta-reasoner to optimize subsequent steps. The selection of the strategy is almost exactly corresponds to the well-studied problem of contextual multi-armed bandits as illustrated in §3. Each strategy can be seen as an arm for the bandit, and the reward of each strategy can be evaluated by the progress of LLM reasoning after applying the strategy. We analogy the process of executing and evaluating each strategy as the act of “pulling” each arm. The overall goal of our meta-reasoner is to find the best arm (i.e., strategy with highest cumulative rewards) with as few pulls as possible. The complete algorithm of Meta-Reasoner is appended in Algorithm 1.
4.1 Chain-of-Thought (CoT) Generation
In the first step, the LLM generates a reasoning step to extend its CoT reasoning based on the user query. Starting from its reasoning history $C_{t-1}$ and the guidance $G_{t-1}$ provided by the meta-reasoner in the previous round, the LLM $M$ produces a new reasoning step $s_{t}$. This step is appended to the current CoT, forming $C_{t}=C_{t-1}\cup\{s_{t}\}$. By maintaining the full reasoning trajectory at each round, the model establishes a coherent foundation for evaluation and further refinement. This approach resembles the long-term reasoning demonstrated in models like o1, which generate extended CoTs. However, this reasoning process often resembles “trial-and-error”, potentially incurring unnecessary inference costs on unproductive paths. Moreover, due to its sequential nature, the process is prone to becoming trapped in suboptimal solutions.
4.2 Progress Reporting
After updating the CoT, we summarize the reasoning history $C_{t}$ into a concise progress report $P_{t}$. This summary captures key aspects of the reasoning trajectory, including progress toward the task goal, reasoning consistency, and significant updates thus far. The summarization function $f$ abstracts the detailed CoT into a simpler, more focused representation. This step is designed to be both computationally efficient and informative, enabling the meta-reasoner to evaluate high-level progress without being overwhelmed by the granular details of every reasoning step. While this may be viewed as an engineering heuristic, we observe that including essential information in the prompt unlocks the LLM’s capacity for “higher-order” thinking. Consequently, the LLM tends to generate more insightful and critical strategies, which are particularly valuable for complex reasoning tasks.
@@ 4.2 Progress Reporting
Table 1: Demonstration: Contextual bandit pair (i.e., diagnosis of the progress report (context) and the corresponding strategy (bandit)) for guiding the LLM’s reasoning process. Marked rows are some of the unique strategies generated by Dynamic Contextual Bandits.
4.3 Meta-reasoner Strategy Generation

### v4

@@
Abstract.
Large Language Models (LLMs) struggle with high computational time and error propagation during inference time, especially for complex tasks like math, puzzles, or coding requiring multi-step thinking. While existing reasoning models with chain-of-thoughts (CoT) can enable LLMs to do step-wise analysis and reflection, they often face the issue of wasting computation on less productive solutions and fail to make progress during inference time. In this paper, we propose Meta-Reasoner, a new framework to enable LLMs “Think about how to think”, i.e., optimize the inference compute by adjusting strategies on how to reason during inference time. Inspired by dual-process theory, our method decouples the high-level strategy generation (e.g., backtracking, switching approaches, or restarting) from stepwise CoT generation via a lightweight progress report. The strategy module only consider the summarized version from the previous CoTs to propose new strategies accordingly. We employ the contextual multi-armed bandits (CMABs) for this module to iteratively evaluate the previous reasoning states and dynamically adjust the strategy to avoid reasoning get stuck in less productive paths during inference. Evaluations on math problems (e.g., Game-of-24, TheoremQA) and scientific problems (e.g., SciBench) demonstrate that our method improves performance by 9-12% over previous SOTA methods while reducing inference time by 28-35%. This approach also generalizes to other domains like creative writing, demonstrating its versatility for diverse reasoning-intensive problems using LLMs.
Large Language Models, Meta-Learning, Dual-Process Theory, Inference Time Compute
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generation
1. Introduction
@@
Recent models incentivize stepwise chain-of-thought (CoT) reasoning (e.g., o1 from OpenAI (OpenAI et al., 2024), and r1 from Deepseek (DeepSeek-AI et al., 2025)) have achieve remarkable performance on complex reasoning tasks, such as math (Patel et al., 2024; Lightman et al., 2023), science (Rein et al., 2024), and logical puzzles (Lei et al., 2024; Yao et al., 2023). By simulating human-like deliberation (Yao et al., 2024; Wei et al., 2022), these methods allow large language models (LLMs) to decompose problems into smaller subproblems, test hypotheses, reflect on intermediate results, or iteratively refine their solutions (Cao et al., 2024). This extended reasoning process enables LLMs to achieve progressive improvement during inference time before producing a final answer (Chenghao Yang, 2024).
Despite these advances, these multi-step reasoning paradigm remains challenged by its trial-and-error nature: models generate numerous candidate reasoning paths, explore each of them and discard the error ones and gradually converge on final solution. While this facilitates exploration of diverse strategies, it suffers from two critical limitations: (1) computational inefficiency, where models waste resources exploring dead-end paths during inference (Snell et al., 2024; Manvi et al., 2025); and (2) error propagation, where early mistakes accumulate and compromise subsequent steps (Lei et al., 2024; Yao et al., 2023; Gandhi et al., 2024). While models like o1 can perform limited backtracking, such as reevaluating individual steps or self-verification to revise their outputs (Gandhi et al., 2024; Li et al., 2025). They still lack a systematic mechanism to holistically revise or redirect their reasoning. This inability to dynamically adapt their strategy can lead to the model becoming stuck on unproductive lines of thought or inefficiently exploring the problem space without recognizing a strategic pivot is necessary (Gao et al., 2024). Additionally, as LLMs are tackling more challenging problems which require longer reasoning paths, these limitations will be critical bottleneck for the LLM development. A critical challenge, therefore, is to enable LLMs to manage their reasoning budget more effectively, i.e., prioritizing promising directions while adapting or discarding ineffective strategies during inference time. Addressing this requires a novel approach to provide adaptive oversight from the reasoning process.
In this paper, we propose Meta-Reasoner, a specialized module that operates alongside the LLM to dynamically optimize its reasoning strategies during inference time. Acting as a high-level advisor, Meta-Reasoner dynamically evaluates the reasoning process of the LLM and provides strategic guidance or redirection when progress stalls. Inspired by dual-process theory (Didolkar et al., 2024), Meta-Reasoner decouples high-level strategies from detailed stepwise CoT generation from LLM via a lightweight progress report module. The strategy generation module only consider the summarized version from the previous CoTs to propose new strategies accordingly. The system operates iteratively: (1) the LLM generates partial chain-of-thought reasoning steps (§4.1) and a summarizes the reasoning states into a lightweight progress report (§4.2). (2) Meta-Reasoner then reviews this report and using contextual multi-armed bandits (CMABs) (§4.3) to generate high-level strategies (e.g., backtrack to step 3, switch decomposition approach, or restart with new axioms). (3) The LLM then executes refined steps guided by this feedback. This setup allows Meta-Reasoner to concentrate on overall strategy rather than getting involved in the granular details of the LLM’s reasoning. Based on our experiments results (§5), we find that our proposed method demonstrate better performance compared to other baselines on complex tasks such as math, puzzles, or logic tasks, and the performance of our method can be generalized to some other domains like creative writing. Overall, the main contributions of this paper are as follows:
@@
  • We propose a meta-reasoning module that provides high-level strategy and guidance for LLM reasoning, to prevent it from getting stuck on unproductive solution paths during inference.
@@
  • The framework requires no task-specific tuning and generalize to math, science, puzzles, and other creative problems require complex reasoning.
  • We empirically show that our Meta-Reasoner improves accuracy and efficiency on challenging benchmarks (e.g., Game of 24, TheoremQA, and SciBench) compared to existing strong baselines. Our results highlight the framework’s potential as a new scalable solution to inference-time reasoning.
2. Related Works
@@
Complex Reasoning in LLMs.
@@
The introduction of CoT reasoning has transformed how LLMs tackle complex problems by decomposing tasks into intermediate steps (Lee et al., 2025). Recent models such as OpenAI’s o1 and o3, and Deepseek’s r1, have achieved state-of-the-art performance across diverse domains by leveraging CoT-like reasoning (Manvi et al., 2025; Li et al., 2025; Kudo et al., 2024; Sui et al., 2025a). However, the inherent sequential dependency in CoT limits robustness: errors in early steps can propagate and degrade overall performance (Snell et al., 2024). Moreover, when confronted with complex reasoning tasks, CoT methods may become trapped in infinite reasoning loops (Lee et al., 2025; Sui et al., 2025b). To address these challenges, we propose Meta-Reasoner, a meta-reasoner that monitors and adapts the reasoning strategy based on the progress of CoT. By integrating mechanisms such as backtracking and self-verification, Meta-Reasoner provides a holistic perspective that evaluates the overall reasoning trajectory to prevent stagnation and reduce wasted computation on unproductive lines of thought.
Backtracking and Self-Verification.
@@
To mitigate the limitations of CoT-like reasoning, recent methods have explored backtracking and self-verification techniques (Yao et al., 2023; Besta et al., 2023; Gandhi et al., 2024). For example, Weng et al. (2023) demonstrate that incorporating a self-verification stage—where the model re-examines its conclusions using the generated chain of thought—significantly improves performance by detecting errors early. Similarly, Ling et al. (2023) propose generating multiple candidate reasoning chains alongside a verifier mechanism that identifies and backtracks on erroneous steps. These approaches extend beyond post-hoc validation by enabling dynamic strategy adjustments during inference (Lightman et al., 2023), thereby limiting error propagation in lengthy reasoning chains and mitigating infinite reasoning loops. Building on these efforts, our Meta-Reasoner framework employs instructions such as (1) restart from scratch with alternative strategies, (2) backtrack to the error point, and (3) continue with targeted suggestions to optimize the reasoning process. Further details on this strategy design are provided in §4.3.
Meta-Cognition & Dual-Process Systems.
@@
From a cognitive science perspective, meta-cognition involves higher-order processes that allow individuals to monitor, evaluate, and adjust their cognitive strategies (Gao et al., 2024; Yoran et al., 2023) during thinking process. This reflective thinking—often characterized as System 2 in dual-process theories (Havrilla et al., 2024)—is vital for tasks requiring careful deliberation and error correction (Didolkar et al., 2024). Drawing on these insights, our Meta-Reasoner framework can be viewed as analogous to dual-process systems: the LLM generates CoT steps akin to System 1, while the Meta-Reasoner provides high-level strategic oversight, analogous to System 2, guiding or redirecting reasoning as needed. This separation of responsibilities balances efficiency with robust problem-solving, allowing the LLM to handle routine inferences and the Meta-Reasoner to intervene for strategic adjustments.
3. Preliminary
In complex reasoning tasks, a key challenge is deciding the best strategy from several valid options. Considering solving a complex math problem: researchers may employ strategies like decomposition, abstraction, heuristic validation, boundary testing or any other methods at any given step. The critical question is which strategy to use and when. This decision making problem aligns well with contextual multi-armed bandits (CMABs) problem (Slivkins, 2019), a well-studied framework for making such choices. The framework is designed to balance exploration (trying new strategies to learn) and exploitation (using strategies known to be effective).
Figure 1. Demonstration: Contextual Multi-armed Bandits.
Intuitive Understanding of CMAB.
@@ 3. Preliminary
Imagine an agent faced with several strategies (called arms), and at each step, it observes some information about the current situation (called the context). Based on this context, the agent picks one strategy to apply and then receives feedback (a reward) indicating how well that strategy performed. The goal of the CMAB problem is to choose strategies over time to maximize the total reward.
As shown in Figure 1, it illustrates a dynamic strategy selection process using the CMAB approach. The controller manages three arms which correspond to three strategies—Strategy 1, Strategy 2, and Strategy 3—each with an associated Strategy Value (Q-value) that reflects its effectiveness. Based on feedback ($r_{t_{b}}$) from problem-solving attempts, the controller updates the probability distribution $\pi_{t_{b}}$ (illustrated as the “strategy choice probability”), which determines the likelihood of selecting each strategy (e.g., Strategy 1 might have a 50% chance, Strategy 2 a 30% chance, and Strategy 3 a 20% chance). The sample process then picks a specific strategy $\alpha_{i}$ (illustrated as the “chosen strategy”) from this distribution to apply to the current problem. The resulting performance feedback is fed back into the controller, which adjusts the Strategy Values and refines $\pi_{t_{b}}$ for the next iteration, optimizing the strategy selection over time. The bar chart visualizes the current distribution of $\pi_{t_{b}}$ across the strategies.
Formalization of CMABs.
@@
(1) $R(T)=\sum_{t=1}^{T}r(s_{t},x_{t}).$
A popular algorithm for this problem is LinUCB (Li et al., 2010), which assumes the expected reward is approximately a linear function of the context. Specifically, for each arm $s$, there is an unknown parameter vector $\theta_{s}$ such that:
(2) $\mathbb{E}[r(s,x_{t})]\approx x_{t}^{\top}\theta_{s}.$
@@
(3) $s_{t}=\arg\max_{s\in\mathcal{S}}\left[x_{t}^{\top}\hat{\theta}_{s}+c\sqrt{x_{t}^{\top}A_{s}^{-1}x_{t}}\right],$
@@
By using context to guide its choices, LinUCB adapts its strategy selection to the current situation, which aligns with our goal in Meta-Reasoner: to dynamically choose the most effective reasoning strategy during inference time. We provide the details about how to leverage these CMABs approaches in our system in Section 4.3.
@@
Figure 2. Workflow of Meta-Reasoner. In each round, LLM produces a new reasoning step to extend its CoT reasoning (§4.1). The CoT is then summarized into a progress report, which provides context for the meta-reasoner (§4.2). Then meta-reasoner uses a contextual multi-armed bandit (either using a fixed contextual bandit or dynamic contextual bandit) to choose a guidance strategy (§4.3). The selected strategy then guides the next reasoning step generation, to enable strategic redirection, error correction, and resource optimization. A reward is then computed from the progress report and used to update the bandit algorithm (§4.3.1). The process repeats until the task is complete or the maximum number of rounds is reached.
Table 1. Demonstration of contextual bandit pairs (i.e., diagnosis of the progress report (context) and the corresponding strategy (bandit)) for guiding the LLM reasoning process. The bottom four rows are some of the unique strategies generated by Dynamic Contextual Bandits illustrated in §4.3.3.
@@
4. Methods
Motivated by the intuition that LLMs should concentrate their computational efforts on more promising reasoning paths during inference time, we explore two key research questions in this paper:
  • (1) How can language models dynamically allocate resources during inference to optimize reasoning and planning?
  • (2) What architectural design enables an effective separation between the reasoning process within the LLM and the meta-level guidance that oversees it?
To address these questions, we propose Meta-Reasoner, which endows LLMs with the ability to “think about how to think”. Our framework supervises the LLM’s reasoning process and dynamically guides the model to focus on more promising reasoning trajectories during inference time. The framework operates iteratively as illustrated in Figure 2. At each round $t$, the reasoning process comprises three steps: (1) CoT generation by the LLM, (2) Progress Reporting to summarize the reasoning progress so far (i.e., this is partly for efficiency, and partly to help the meta-reasoner focus on its main goal of “advising” rather than being distracted by the details in the CoT), and (3) Strategy Generation by the meta-reasoner to optimize subsequent steps. The selection of the strategy is almost exactly corresponds to the well-studied problem of contextual multi-armed bandits as illustrated in §3. Each strategy can be seen as an arm for the bandit, and the reward of each strategy can be evaluated by the progress of LLM reasoning after applying the strategy. We analogy the process of executing and evaluating each strategy as the act of “pulling” each arm. The overall goal of our meta-reasoner is to find the best arm (i.e., strategy with highest cumulative rewards) with as few pulls as possible. The complete algorithm of Meta-Reasoner is appended in Algorithm 1.
4.1. Chain-of-Thought (CoT) Generation
In the first step, the LLM generates a reasoning step $s_{t}$ to extend its CoT based on the user query. This step builds on the previous reasoning history $C_{t-1}$ and incorporates strategic guidance $G_{t-1}$ from the meta-reasoner in the prior round. The new reasoning step is appended to the CoT, forming $C_{t}=C_{t-1}\cup\{s_{t}\}$. By maintaining the complete reasoning trajectory at each iteration, the model maintains a coherent foundation for evaluating the refining the progress of reasoning process. This step resembles the long-term reasoning seen in models like o1 and r1, which generate extended CoTs during inference time. The key difference is that our meta-reasoner model actively intervenes the reasoning process by providing strategical guidance $G_{t-1}$, enabling the model “think about how to think”.
4.2. Progress Reporting
After updating the CoT, we apply a summarization function $f:C_{t}\rightarrow P_{t}$ to distill the most recent $t$ steps of the CoT into a concise progress report $P_{t}$. This report encapsulates critical elements of the reasoning trajectory, including the degree of progress toward the task objective, the logical consistency of the reasoning steps, and significant milestones or updates achieved. The function $f(\cdot)$ is designed to be computationally efficient, producing a compact summary that enables the meta-reasoner to assess high-level progress without processing the granular details of each step in $C_{t}$. While this summarization is heuristic in nature, we observe that it may trigger LLMs’ capacity for higher-order reasoning by including only essential information in $P_{t}$, which informs the meta-reasoner’s strategic guidance $G_{t}$. Empirically, this approach fosters more insightful and adaptive strategies, as the meta-reasoner can focus on critical patterns rather than exhaustive step-by-step details. We provide the detailed experiments in §5.3.
4.3. Meta-reasoner Strategy Generation